EX 4-3 Single plantwide factory overhead rate

Salty Sensations Snacks Company manufactures three types of snack foods: tortilla chips,

potato chips, and pretzels. The company has budgeted the following costs for the upcoming

period:

Factory depreciation $ 31,360

Indirect labor 78,400

Factory electricity 7,840

Indirect materials 35,400

Selling expenses 25,000

Administrative expenses 18,000

Total costs $196,000

Factory overhead is allocated to the three products on the basis of processing hours. The

products had the following production budget and processing hours per case:

Budgeted Processing Hours

  Volume(Cases) per Case   

Tortilla chips 4,000 0.20

Potato chips 5,000 0.15

Pretzels     2,500     0.40

Total 11,500

Instructions

Determine the single plantwide factory overhead rate.

Use the factory overhead rate is to determine the amount of total and per-case

factory overhead allocated to each of the three products under generally accepted accounting

principles.
Answer #1

Question A

Single Plantwide Overhead Rate = Total Overhead Costs / Total Processing Hours

Total Overhead Costs = $ 153,000

Total Processing Hours = 2,550 Hours

Single Plantwide Overhead Rate = 153,000 / 2,550

Single Plantwide Overhead Rate = $ 60 per Processing Hour

Note

Total Overhead Cost = Total Costs - Administrative Expenses - Selling Expense

Total Overhead Cost = 196,000 - 25,000 - 18,000

Total Overhead Cost = $ 153,000

Calculation of Total Processing Hours

Tortilla Chips = 4,000 Cases * 0.20 Hours per Case = 800 Hours

Potato Chips = 5,000 Cases * 0.15 Hours per Case = 750 Hours

Pretzels = 2,500 Cases * 0.40 Hours per Case = 1,000 Hours

Total Hours = 800 + 750 + 1,000 = 2,550 Processing Hours

Question B

Calculation of Overhead Total and Per Unit

Tortilla Chips = 800 Hours * $ 60 per Hour = $ 48,000

Overhead per Unit = 48,000 / 4,000 Units

Overhead per Unit = $ 12 per Unit

Potato Chips = 750 * $ 60 per Hour = $ 45,000

Overhead Per Unit = 45,000 / 5,000 Units = $ 9 per Unit

Pretzels = 1,000 Hours * $ 60 per Hour = $ 60,000

Overhead per Unit = 60,000 / 2,500 Units = $ 24 per Unit

Per Unit means Per Case.